OpenStudy (kj4uts):

Which of the following are right triangles? Check all that apply. A. A triangle with side lengths 6 inches, 8 inches, 10 inches B. A triangle with side lengths 8, 15, 17 C. A triangle with side lengths 5, 12, 13 D. A triangle with side lengths 4, 5, 6 Please explain answer. Is there a way to solve this or ratios to remember? Thank you!

OpenStudy (haseeb96):

according to the definition of Right triangle ''sum of two sides equal to the other one side'' okay ?

OpenStudy (haseeb96):

they are all right triangles i will do only one then u will do one by one as i do 

OpenStudy (haseeb96):

now apply Pythagoras theorem that is (hypotenuse)^2 = ( perpendicular)^2 +(base)^2  (6)^2 = (5)^2 + (4)^2 36 =25 +16 36 =36 according to the definition it is right triangle okay now do the remaining questions like this